# AI in Supply Chain Risk Assessment: A Systematic Literature Review and Bibliometric Analysis

### v2

@@ 4.11 Journal Publication Trends
Table 3: Publication Trends by Journal (2015-2023)
Source Title 2015 2016 2017 2018 2019 2020 2021 2022 2023 Total
SUSTAINABILITY
1
2
4
5
10
3
25
INTERNATIONAL JOURNAL OF PRODUCTION RESEARCH
1
1
5
1
9
3
20
ANNALS OF OPERATIONS RESEARCH
5
5
5
15
COMPUTERS INDUSTRIAL ENGINEERING
2
1
1
3
3
3
13
IEEE ACCESS
2
1
4
2
9
EXPERT SYSTEMS WITH APPLICATIONS
1
1
1
1
2
2
8
IEEE TRANSACTIONS ON ENGINEERING MANAGEMENT
2
5
1
8
INDUSTRIAL MANAGEMENT DATA SYSTEMS
1
4
1
2
8
ENGINEERING APPLICATIONS OF ARTIFICIAL INTELLIGENCE
1
1
2
2
1
7
JOURNAL OF CLEANER PRODUCTION
2
5
7
WIRELESS COMMUNICATIONS MOBILE COMPUTING
3
4
7
EUROPEAN JOURNAL OF OPERATIONAL RESEARCH
1
1
3
1
6
INTERNATIONAL JOURNAL OF PRODUCTION ECONOMICS
1
1
1
1
1
1
6
JOURNAL OF INTELLIGENT FUZZY SYSTEMS
1
2
3
6
MATHEMATICAL PROBLEMS IN ENGINEERING
1
1
1
3
6
COMPUTATIONAL INTELLIGENCE AND NEUROSCIENCE
1
4
5
MOBILE INFORMATION SYSTEMS
1
4
5
NEURAL COMPUTING APPLICATIONS
1
2
1
1
5
TRANSPORTATION RESEARCH PART E-LOGISTICS AND TRANSPORTATION REVIEW
1
2
2
5
COMPUTERS IN INDUSTRY
1
2
1
4
Table 3 presents a comprehensive breakdown of published papers in various source titles from 2015 to 2023. These source titles encompass a range of journals and publications contributing to the field of SCRA. The table clearly shows the publication trends, indicating how many papers were published in each source title each year and their cumulative total. It provides valuable insights into the distribution of research output and the source titles that have been active in this area.
Through these figures and tables, our systematic literature review is empowered to offer an in-depth analysis of the SCRA landscape and provide valuable insights to both the academic and professional communities in this dynamic field.
5 AI Techniques for Supply Chain Risk Assessment
In recent years, the application of Machine Learning (ML) methods in Supply Chain Risk Assessment (SCRA) has garnered substantial attention due to its potential to enhance decision-making processes and mitigate operational risks. This review paper explores various ML techniques applied in SCRA, highlighting their methodologies, key findings, and contributions.
Han and Zhang, (2021) employed Factor Analysis and a Backpropagation Neural Network (BPNN) to identify high and extremely high-risk incentives in sustainable supply chains, offering valuable insights for risk assessment. Jianying et al., (2021) optimized a BPNN using Genetic Algorithm (GA) and Particle Swarm Optimization (PSO), enhancing its predictive accuracy, which can be instrumental in the fresh grape industry’s sustainable supply chain management.
Belhadi et al., 2021b utilized a hybrid Ensemble Machine Learning (EML) approach involving the Gama Test, Rotation Forest (RotF), and Logit Boosting (LB) algorithms, achieving excellent performance with the RotF-LB model, contributing to credit risk assessment in the context of Agriculture 4.0. Wu et al., 2022c improved credit risk prediction in agricultural SCF using a GA-BPNN Credit Risk Model, attaining a risk prediction accuracy above 0.92. Zhang et al., (2015) compared different credit risk assessment models for SCF, highlighting the superiority of the Support Vector Machine (SVM) model with an accuracy of 93.65%. Lei et al., (2023) demonstrated the effectiveness of the Support Vector Machine and Slime Mould Algorithm (SVM-SMA) model for financial risk assessment, with results showing Precision: 85.38%, F-score: 63%, and TNR: 72%.
Yin et al., (2022) introduced a Convolutional Neural Network (CNN) model for early warning of SCF risks, achieving an optimal accuracy at 200 iterations and a comprehensive accuracy of 94.7%. Chen et al., (2021) employed a Long Short-Term Memory (LSTM) network model to forecast oil import risk, achieving superior forecasting accuracy compared to other models, offering a more effective risk assessment solution for oil import scenarios. Janjua et al., (2023) developed a real-time social media data assessment tool for SCRM with the Bi-Directional Long Short-Term Memory Conditional Random Field (Bi-LSTM CRF) model, resulting in improved precision, recall, and F1 scores compared to baseline models. Yao et al., (2022) showcased the power of combining multiple ranking information and an ensemble feature selection method for enterprise credit risk assessment, obtaining the best prediction results.Aboutorab et al., (2022) demonstrated the application of Reinforcement Learning (RL) in proactive risk identification, achieving high accuracy in identifying various risks. Kosasih et al., (2022) presented a tensorization-based Neuro-Symbolic model that outperformed existing models and unveiled hidden supply chain risks.
Wang et al., 2022a introduced a novel cost-sensitive learning Random Forest model (CSL-RF) equipped with an imbalance sampling strategy to predict SMEs credit risk in SCF, achieving exceptional accuracy (97.%) and precision (1.00). Wang et al., (2021) addressed the small sample size issue in SME credit risk forecasting by proposing an Adaptive Heterogeneous Multiview Graph Learning (AH-MGL) method, which effectively improved interpretability and demonstrated robustness with an accuracy of 87.44%. Zhu et al., (2016) introduced the Random Subspace-Real AdaBoost (RS-RAB) method, showcasing its effectiveness in credit risk prediction, and emphasized the performance of ensemble methods against individual machine learning techniques, shedding light on their predictive capabilities. Bassiouni et al., (2023) presented a unique case study utilizing Deep Learning (DL) methodologies, such as Recurrent Neural Networks (RNN) and Temporal Convolution Neural Networks (TCN), to predict the exportability of shipments during the COVID-19 pandemic, significantly improving classification accuracy.
Zhu et al., (2019) introduced an enhanced hybrid Ensemble Machine Learning (EML) approach (RS-MultiBoosting) for forecasting SMEs’ credit risk, demonstrating its superiority, especially with small datasets. Athaudage et al., (2022) proposed a Random Forest (RF) regression model for crude oil price analysis during disease outbreaks, offering improved forecasting capabilities with high accuracy. Liu and Huang, (2020) combined AdaBoosting and SVM with Adaptive Mutation Particle Swarm Optimization (AM-PSO) and noise reduction techniques for risk assessment in SCF, resulting in higher credit assessment accuracy. Zhu et al., (2017) analyzed various EML methods, highlighting the effectiveness of the Random Subspace Boosting (RS-Boosting) approach in SME credit risk prediction.
Fu et al., (2022) explored the application of BPNN, SVM, and GA in risk assessment. The results highlighted the superiority of the BP-GA model in terms of classification accuracy, emphasizing the importance of innovative approaches to risk assessment. Zhang, (2022) introduced the Fuzzy Support Vector Machine Prediction Model (FSVM) for enterprise SCRA, demonstrating higher accuracy compared to traditional models, including SVM and Decision Trees (DT), and the neural network model Temporal-Spatial Fuzzy Neural Network (T-SFNN). This research addressed the need for company-focused SCRM solutions, especially in the context of macroeconomic environments and epidemics. Pan and Miao, (2023) presented a DL-based BPNN model for SCRA, showcasing exceptional generalization ability and highlighting its potential as an effective risk assessment method. Yang et al., (2021) introduced the Least Absolute Shrinkage and Selection Operator Logistic Regression (Lasso-Logistic) model, effectively overcoming multicollinearity and offering the best prediction accuracy and discrimination ability.
Zhang et al., (2022) introduced the DeepRisk model, emphasizing the importance of utilizing both static and dynamic data in credit risk prediction for SMEs in SCF. This model outperformed baseline methods in various credit risk prediction metrics, emphasizing the value of dynamic financing behavioral data. Wang, (2021) integrated Blockchain and Fuzzy Neural Network algorithms to study credit risk in SME financing within the SCF context. This approach significantly contributes to credit risk understanding and management within SCF. Li, (2022) applied BPNN and LR analysis to predict investment risk in SCM, demonstrating high prediction accuracy, which can be instrumental in improving SCM and the sustainable development of enterprises. Bodendorf et al., (2023) introduced a DL model that exhibited high predictive performance and enhanced interpretability for managers, answering the ”why” question. This approach represented an extension to SCRM research.
Liu et al., (2022) utilized a hybrid model chain, incorporating eXtreme Gradient Boosting (XGBoost), Synthetic Minority Oversampling Technique for Nominal and Continuous (SMOTENC), and Random Forest (RF), to identify and control credit risk in financial institutions. This approach contributed to developing integrated models that address interpretability and accuracy issues. Rao and Li, (2022) implemented various ML algorithms, including Logistic Regression (LR), Decision Trees (DT), and integrated Logistic Regression-Random Subspace (LR-RS) and Decision Trees-Random Subspace (DT-RS) methods to enhance risk assessment and behavior prediction in blockchain SCF. The Logistic Regression-Random Subspace (LR-RS) model demonstrated practicality and superiority in credit risk assessment. Zheng et al., (2023) introduced a Federated Learning (FL) method to address supply chain risk prediction. FL enabled supply chain members with smaller and imbalanced datasets to leverage collective information, enhancing predictive performance. It outperformed other algorithms, demonstrating the potential of collective learning in supply chain risk prediction.
Handfield et al., (2020) presented a method for predicting factory risks in Low-Cost Countries (LCCs) with a five-year planning horizon, enabling proactive risk assessment before sourcing decisions. This approach addressed the geographic-specific risks associated with supplier factories and informed sourcing allocation decisions. Wong et al., (2022) highlighted the significance of AI-driven risk management in enhancing Supply Chain Agility and Re-engineering Capabilities (RP). This study explored the potential of artificial intelligence in SME risk assessment and its impact on business continuity. Wang and Song, (2022) utilized the Interval Type 2 Fuzzy Neural Network (IT2FNN) method, effectively addressing periodic deception and exhibiting a high comprehensive accuracy rate. The research combined fuzzy logic systems and artificial neural networks to assess e-commerce credit risk.
Nezamoddini et al., (2020) introduced a novel GA integrated with an Artificial Neural Network (ANN) to optimize supply chain decision-making, emphasizing the robustness and profitability improvements the proposed technique offers. Zhang, (2022) presented an SVM based on a Fuzzy Model (FSVM) for enterprise SCRA. This method’s accuracy surpassed traditional models, contributing to the effectiveness of complex enterprise SCRA. Xia et al., (2023) explored ML methods for manufacturing SCF credit risk assessment. RF emerged as the best-performing model, and the research focused on analyzing the unique environment of Small and Medium-Sized Manufacturing Enterprises (SMMEs) in China. Zhao and Li, (2022) examined the use of SVM and BPNN algorithms for measuring risk in SCF, aiming to address the financing dilemma of SMEs.
Li and Fu, (2022) highlighted the superiority of the Principal Component Analysis-Genetic Algorithm-Support Vector Machine (PCA-GA-SVM) model for credit risk prediction in the specific context of SCF accounts receivable mode. The model exhibited better performance in terms of accuracy and error rates. Duan et al., (2021) employed a BPNN optimized by the GA model. This approach offered improved fitting effects, higher prediction precision, and faster convergence than conventional models. Zhang et al., (2021) introduced the Firefly Algorithm Support Vector Machine (FA-SVM) for financial credit risk assessment in supply chains. The FA-SVM enhanced classification efficiency and reduced error rates, making it more practical.
Luo et al., (2022) used SVM with improved optimization techniques, such as Dynamic Mutation Particle Swarm Optimization (DM-PSO), to optimize parameters. The integration of AdaBoost as a weak classifier resulted in an integrated model with superior performance. Dang et al., (2022) constructed a BPNN and incorporated blockchain technology for credit risk evaluation in SCF. The research demonstrated the effectiveness of deep learning and blockchain in supply chain financial prediction and management. Hosseini and Barker, (2016) proposed a Bayesian Network (BN) model to evaluate and select suppliers, considering criteria falling into primary, green, and resilience categories. This model quantified supplier resilience in absorptive, adaptive, and restorative capacities, addressing the importance of resilience in the supplier selection process. Gao et al., (2022) explored various ML models, including Autoregressive Integrated Moving Average (ARIMA) and Multilayer Perceptron-Long Short-Term Memory (MLP-LSTM), for commodity demand prediction. The research highlighted the effectiveness of the Autoregressive Mixture Density Network (AR-MDN) commodity prediction model and the Improved Particle Swarm Optimization (IPSO) algorithm.
Wei, (2022) introduced a Machine Learning-based Linear Regression Algorithm (ML-LRA) for credit risk assessment. This method efficiently reduced the risk of supplier credit risk assessment, contributing to the field of supply chain management. Deiva Ganesh and Kalpana, (2022) employed text mining to analyze live supply chain-related information from social media platforms, emphasizing the importance of real-time risk identification. Zhang et al., (2023) presented a Time-Decayed Long Short-Term Memory (TD-LSTM) algorithm for monitoring and interpolating missing data in irregular time series. This algorithmic model enhanced the predictability of datasets with missing data. Baryannis et al., 2019a proposed a Generic Data-Driven Risk Prediction Framework (GDDRPF) for SCRM. The framework emphasized the synergy between Artificial Intelligence (AI) and supply chain experts, considering the trade-off between interpretability and prediction performance.
Collectively, these papers contribute to the evolving field of supply chain risk assessment and management, leveraging various machine learning, optimization, and data-driven approaches to enhance prediction, decision-making, and resilience within supply chains.

### v3

@@ 4.11 Journal Publication Trends
Table 4 presents a comprehensive breakdown of published papers in various source titles from 2015 to 2024. These source titles encompass a range of journals and publications contributing to the field of SCRA. The table clearly shows the publication trends, indicating how many papers were published in each source title each year and their cumulative total. It provides valuable insights into the distribution of research output and the source titles that have been active in this area.
5 Framework for SCRA using AI
Figure 14: Framework for SCRA using AI.
Our proposed framework for AI-driven SCRA provides a comprehensive roadmap for SC professionals to integrate AI techniques effectively. Figure 14 visually shows the sequential steps involved, starting from defining key risk indicators to implementing preventive measures. By following these steps, practitioners can systematically define risk indicators, source and pre-process data, select and train models, and interpret results to facilitate informed decision-making. The framework is enriched with insights from existing studies on AI and ML applications in SC risk, presenting best practices, key considerations, and real-world challenges and opportunities identified in the literature.
5.1 Defining Key Risk Indicators
Our first step is identifying the key input and output variables that drive SCRA. Input variables include transactional data, supplier metrics, and external factors like political or environmental conditions, while output variables classify risks into severity levels such as “low,” “medium,” and “high” to align with organizational decision-making priorities. Defining these variables is crucial, as they represent the foundation upon which AI models assess risk. Wu, (2024) emphasize capturing operational and risk-related data for comprehensive risk assessments, with Xu et al., (2024) using historical sales and market trends for disruption prediction. Including financial and non-financial indicators, as highlighted by Wang et al., (2021), enhances the assessment of risk exposure. Practitioners should ensure these variables align with their industry’s risk characteristics to produce actionable and context-relevant insights.
5.2 Data Collection
Effective risk assessment relies on both internal and external data sources to capture a comprehensive picture of potential risks. Internal data from ERP systems, procurement records, and inventory management systems provides insights into operational processes. External data, such as market indicators, social media sentiment, and climate reports, are critical for Deiva Ganesh and Kalpana, (2022) used real-time data from social media, that is, Twitter, as the input source. ERP systems, supplier systems, and logistics systems, alongside external sources like social media and weather forecasts. Such data covers both structured (e.g., numerical data) and unstructured formats (e.g., text or multimedia) to enhance risk detection. Deiva Ganesh and Kalpana, (2022) used real-time data from social media, that is, Twitter, as the input source. The author used Twitter streaming API as a primary source for data extraction. Kosasih et al., (2022) utilized external platforms like Marklines for automotive SC data and Achilles Information Ltd for energy sector insights. Pan and Miao, (2023) collected data through surveys from industry practitioners to capture qualitative and contextual insights about SCR.Including such data enhances the model’s ability to preemptively detect risks like demand fluctuations or reputational issues, underscoring the value of a multimodal data approach.
5.3 Data Preprocessing
Preprocessing is a vital step. Raw data must undergo preprocessing to enhance model accuracy and efficiency. This step involves cleaning data to remove noise, normalizing values, and encoding categorical variables. Dimensionality reduction techniques like principal component analysis (PCA) applied by Luo et al., (2022) and imputation methods like residual mean replacement applied by Zhang et al., (2021) address challenges such as high-dimensional datasets and missing values. Time-series data, common in SC applications, requires additional adjustments like seasonality corrections. Techniques like Isolation Forest and Local Outlier Factor (LOF) detect anomalies, enhancing prediction reliability. SPSS statistical preprocessing methods have also been employed effectively in by Dang et al., (2022).
5.4 Model Selection
Practitioners have a range of model choices, each suited to different data characteristics and objectives in SCRA. Classical models like Support Vector Machine (SVM), Decision Tree (DT), and Logistic Regression(LR) are ideal for well-structured, low-dimensional datasets, where the data relationships are more straightforward and computational efficiency is prioritized. DL models, such as Convolutional Neural Networks (CNNs) and Long Short-Term Memory (LSTM) networks, are valuable for handling complex, high-dimensional data, including time-series demand data, where capturing temporal dependencies is essential. Hybrid models, which combine conventional machine learning (CML) and DL approaches, balance interpretability and accuracy, catering to scenarios where insights into model decisions are as important as predictive performance. Ensemble methods like Random Forest(RF) and XGBoost aggregate outputs from multiple models to enhance predictive accuracy and robustness. Additional research emphasizes that model selection should consider factors such as data availability, computational resources, and the necessity for model interpretability. Highly interpretable models are particularly important in regulated industries, such as finance, where transparency is critical. Hybrid and ensemble models, increasingly popular in recent literature, often effectively compromise predictive power and interpretability, making them suitable for complex SC risk scenarios.
5.5 Model Training and Hyperparameter Tuning
Model training typically involves cross-validation to improve model robustness. Tuning parameters like learning rate, depth, and number of estimators (in ensemble methods) is crucial to optimize performance. Tailored hyperparameter tuning can significantly enhance model performance in SC applications. Many researchers highlight the importance of automated tuning techniques, such as grid search, random search, and Bayesian optimization, to streamline model refinement and ensure that the final model is well-suited for specific risk scenarios. Wu et al., 2022b emphasized iterative optimization, while Zhang et al., (2022) addressed data imbalance with SMOTENC. Practitioners may also consider transfer learning approaches to leverage knowledge from similar datasets, especially in situations with limited labeled data.
5.6 Prediction
Once trained, the models generate predictions in the form of classifications (e.g., “high risk” vs. “low risk”), severity scores, or probabilities, which guide risk mitigation strategies. Baryannis et al., 2019b highlighted the role of predictive analytics in shielding SCs from disruptions. These predictions should ideally integrate into real-time dashboards to allow continuous monitoring and provide quick insights into risk fluctuations. This step facilitates proactive risk management, enabling organizations to adjust strategies promptly based on predicted risk levels and preventing disruptions before they escalate.
5.7 Interpretation of Results
Interpretability makes AI-driven predictions actionable, especially for stakeholders without a technical background. XAI, such as SHAP values or LIME, helps practitioners understand which variables most influence predictions. According to Olan et al., (2024), techniques like SHAP and LIME enhance decision-making by elucidating the rationale behind ML predictions, enabling better interpretability and fostering trust in SCRA. Interpretation tools also validate results with domain experts, ensuring the model’s output aligns with real-world SC conditions and providing a basis for actionable policy recommendations.
5.8 Preventive Measure
Based on the risk predictions and insights, practitioners can establish preventive strategies to mitigate potential SC disruptions. Common actions include adjusting inventory buffers, diversifying suppliers, or modifying production schedules. Preventive measures should be continuously refined based on AI predictions and risk evaluations. Incorporating feedback loops that allow practitioners to assess the impact of preventive measures on risk reduction is a good practice. Such loops can help organizations refine their approach over time, ensuring that the AI-driven risk assessment framework remains adaptable to changing conditions.
Table 5: Methods used and the key findings from the review articles.
Reference(s) Methods/Models Key Findings
Katsaliaki et al., (2022) Tensorization-based neural symbolic model Outperformed SNLP and GNN-LP in predicting relationships (AUC: 0.82, AP: 0.72).
Wang et al., (2022) Cost-sensitive learning to a RF (CSL-RF) model Improved prediction accuracy (Accuracy: 97.22%, AUC: 0.9850).
Wang et al., (2021) Adaptive heterogeneous multiview graph learning Improved forecasting accuracy (Accuracy: 87.44%).
Jianying et al., (2021) Optimized Backpropagation Neural Network (BPNN) and Particle Swarm Optimization-Backpropagation (PSO-BP) Outperformed the single Backpropagation (BP) (MAE: 2.1022, $R^{2}$: 0.932).
Chen et al., (2021) LSTM for forecasting LSTM model suitable for risk prediction (MAE: 1.037 - 1.52).
Zhu et al., (2016) RS-RAB Best prediction performance (Accuracy: 86.74%).
Bassiouni et al., (2023) Various DL approaches DL models improved accuracy by 14.46% (Softmax: 76.32%, SVM: 85.23%).
Zhu et al., (2019) RS-MultiBoosting superior performance of the RS-MultiBoosting ensemble method in forecasting
Athaudage et al., (2022) RF regression RF suitable for price prediction (MAE: 1.077, RMSE: 1.56, MAPE: 4.03%).
Liu and Huang, (2020) AdaBoosted SVM (EN-AdaPSVM) Outperforms other SVM models in noise data (Total False Rate: 11.76% (EN-AdaPSVM), 13.41% (AdaPSVM), 32.94% (CSVM)).
Zhu et al., (2017) Ensemble methods (RS-boosting and multi-boosting) RS-boosting outperforms other methods (Average Accuracy: 85.41% (RS-boosting), 84.08% (multi-boosting), 74.80% (boosting)).
Fu et al., (2022) BPNN, SVM, GA BP-GA model outperforms BP and SVM (Accuracy: BP-GA 97.19%, BP 82.33%, SVM 95.31%).
Zhang, (2022) Multimodal DL (DeepRisk) DeepRisk outperforms baseline methods.
Yang et al., (2021) Lasso-logistic model Lasso-logistic model has best prediction accuracy (Prediction Accuracy: 96.5%, Type II Error: 0.037).
Zhang et al., (2022) Fuzzy SVM (FSVM) FSVM outperform traditional SVM and DT (Higher accuracy in training and test sets).
Wang, (2021) Blockchain and fuzzy neural networks Average Accuracy: 84.11%, 84.66%(Independent Variables Set Vs).
Li, (2022) BPNN, LR BPNN and LR provide accurate predictions (Prediction Accuracy: 92.8%).
Bodendorf et al., (2023) DL model with causal inference High predictive performance and causal analysis.
Liu et al., (2022) Hybrid model (XGBoost–SMOTENC–RF) Development of a hybrid model chain, incorporating XGBoost, SMOTENC, and RF models (overall accuracy rate 91.67% and the average percentage error is 6.39%).
Rao and Li, (2022) LR, DT, integrated Logistic-RS Integrated RS improves prediction performance (Accuracy of Logistic: 0.831, Accuracy of Logistic-RS: 0.871, Accuracy of DT: 0.772, Accuracy of DT-RS: 0.795).
Handfield et al., (2020) Newsfeed analysis Predicted factory risks for a five-year planning horizon.
Pan and Miao, (2023) DL BPNN BPNN has high accuracy and low relative error (The maximum relative error calculated by the AHP is 57.41%, and the minimum is 5.88%, while the maximum relative error obtained by the BPNN is only 0.00210526%).
Wong et al., (2022) PLS-SEM, ANN AI-based models (AIRM) are significant determinants of RP and SCA.
Wang and Song, (2022) Interval Type 2 Fuzzy Neural Network (IT2FNN) IT2FNN punishes periodic deception, high accuracy (Accuracy 89%).
Nezamoddini et al., (2020) GA with NN Improved profits, lower inventory levels.
Xia et al., (2023) SVM, RF, MLP, LR RF showed best accuracy and low type I error.
Zhao and Li, (2022) SVM and BPNN Effective risk measurement, low relative error.
Li and Fu, (2022) PCA-GA-SVM PCA-GA-SVM outperforms SVM and GA-SVM, high accuracy.
Duan et al., (2021) BPNN optimized by GA Better fitting effect, higher prediction precision, and higher convergence speed.
Zhang et al., (2021) SVM optimized by firefly algorithm (FA-SVM) Improved classification efficiency and reduced error rates.
Luo et al., (2022) SVM, Particle Swarm Optimization, AdaBoost Improved classification effect, highest accuracy of 96.13%.
Dang et al., (2022) DL technology (BPNN) and blockchain Effective application of DL and blockchain.
Hosseini and Barker, (2016) Bayesian network (BN) BN allows disruption and improvement analysis.
Gao et al., (2022) Autoregressive integrated (AR), Mixture Density Networks (MDN) model, Autoregressive Integrated Moving Average (ARIMA) model, Multilayer Perceptron-Long short term memory (MLP-LSTM) model, Particle Swarm Optimization (PSO) algorithm and the improved PSO (IPSO) algorithm AR-MDN outperforms other models, IPSO has strong performance.
Wei, (2022) ML-based linear regression algorithm (ML-LRA) ML-LRA reduces supplier credit risk.
Zheng et al., (2023) Federated learning with dimensionality reduction (PCA) Collective risk prediction benefits organizations with small datasets (Maximum Difference: 0.0670 (FL vs. Local Learning), 0.0634 (FL vs. Centralized Learning).
Deiva Ganesh and Kalpana, (2022) Text-mining Use of social media data for contemporary risk understanding.
Zhang et al., (2023) Time-decayed long short-term memory (TD-LSTM), Deep Neural Network Improved predictability through TD-LSTM interpolation.
Baryannis et al., 2019a SVM and tree learning Framework offer good performance with interpretable and black-box methods (SVM Accuracy: 0.943, DT Accuracy: 0.950, Restricted DT (RDT) Accuracy: 0.916).
Han and Zhang, (2021) BPNN model BPNN effectively predicted high to extremely high risks in a food production company’s SC.
Belhadi et al., 2021b Hybrid EML approach: Gama Test (GT), Rotation Forest algorithm (RotF), Loogit Boosting algorithm (LB) Good performance of RotF-LB model (Best results for RotF-LB: Accuracy 91.72, Precision 92.96, Recall 95.53, F-measure 94.23, AUC 92.89).
Wu et al., 2022a BPNN and GA-BPNN Credit Risk Model Comparing iterations and risk prediction accuracy (GA-BPNN risk prediction accuracy above 0.92).
Lei et al., (2023) Binary image difference common area (BCAoID), SVM-SMA Highlight the effectiveness of the CGOA-SVM-SMA algorithm (Precision: 85.38%, F-score: 63%, TNR: 72%).
Yin et al., (2022) Convolutional neural network model (CNN) Optimal accuracy at 200 iterations (Comprehensive accuracy: 94.7%).
Janjua et al., (2023) Bi-LSTM CRF model Comparing Bi-LSTM CRF with Baseline CRF (the Bi-LSTM CRF model achieved an overall F1-Score of 85%, the Baseline CRF model achieved an overall F1-Score of 80%).
Yao et al., (2022) Hybrid model combining FS-MRI, SVME-AIR, and SMV AUC and KS scores for different parameters (AUC score of 0.8772 when the parameter “num_AIR” was set to 10, KS score of 0.6363 when the parameter “num_AIR” was set to 4).
Zhang et al., (2015) SVM and BPNN SVM model is superior to the BPNN model (BPNN Accuracy 66.67%, SVM model Accuracy 76.67%).
Aboutorab et al., (2022) RL-PRI approach Time efficiency of RL-PRI vs. manual process (Tarif Dispute Accuracy: 0.941463, Recession Accuracy: 0.907937, Sea Level Rise Accuracy: 0.964286, Flood Accuracy: 0.856322).
Wu, (2024) NLP techniques, word embedding models, bag-of-words approach Developed SCRisk measure, significantly correlated with operational buffers and stock return volatility.
Xu et al., (2024) Large language models (e.g., GPT-4, BERT) Improved decision support, optimized inventory management, reduced transportation costs, risk prediction.
Chen et al., (2024) GA-BPNN, Logistic Regression High accuracy in predicting MSME default and credit ratings (Accuracy: 92% for default prediction, 86% for credit rating).
6 Content Analysis of AI Techniques in SCRA
In recent years, the application of ML methods in SCRA has attracted substantial attention due to its potential to enhance decision-making processes and mitigate operational risks. This review paper explores various ML techniques applied in SCRA, highlighting their methodologies, key findings, and contributions.